OpenStudy (anonymous):

solve the equation (3k+4)^2=64

OpenStudy (anonymous):

solve it like this (3k+4)(3k+4)=64

OpenStudy (anonymous):

what s the answer?

OpenStudy (anonymous):

-4,4/3

OpenStudy (anonymous):

Thanks!!!

jhonyy9 (jhonyy9):

(3k+4)2=(3k+4)(3k+4)=64=8 squared than 3k+4=8 3k=4 k= 4/3

jhonyy9 (jhonyy9):

so prove the (-4) than

OpenStudy (anonymous):

f(-4)=(3(-4)+4)^2=64 =(-12+4)^2=64 =(8)^2 =64 64=64 (this proves -4 for you jhonny9) this is a weird question